What are the possible outcomes when a die is rolled?
If a normal die is rolled once, the outcomes are {1, 2, 3, 4, 5, 6}.
What is the probability that a gaussian probability density function will take a value less than -2?
Assuming that you are considering a N(0,1) Gaussian distribution, the answer is approximately 1 in 20. The 0.95%ile (two tailed) occurs at -1.96 and 1.96.
3
Word 1) 'math' has one vowel letter among a total of 4 letters. The probability of
randomly selecting the vowel letter 'a' is P(v) = 1/4.
Word 2) 'jokes' has two vowel letters among a total of 5 letters. The probability of
randomly selecting a vowel letter is P(v) = 2/5.
The probability of randomly selecting a vowel letter from the first word and a vowel
letter from the second word is: P(v1,v2) = 1/4 (2/5) = 2/20 = 1/10 = 0.10 = 10.0%
50-50
If you have 5 colors on a spinner what is the probability it would land on red?
Assuming that the colors are balanced, the probability is 1 in 5.
Time-Probability Graph
Denote by mu the mean of X. Given
P(X>5.52)=0.305
i.e.,
P((X-mu)/4> (5.52-mu)/4)=0.305
So, letting Z=(X-mu)/4. Then Z~N(0,1). Hence, the above equality is
P(Z>(5.52-mu)/4)=0.305. So,
(5.52-mu)/4=0.51
So,
mu=5.52-4*0.51=3.48
So, the best answer is 3.5
A card is drawn from a pack of 52 cardsWhat is the probability that the card is an ace or a king?
513
0 (there is no chance it will be short since tall is dominant over short).
Hope this helps!
- Biology Student
Since the numbers (number of dots) on normal dice only go from 1 - 6, then I would think the probability of getting the number 100 is zero.
There are 6 different doubles that can be rolled. (They are 1 and 1, 2 and 2, etc.) You have a 6/36 chance of rolling doubles. The fraction 6/36 simplifies (reduces) to 1/6. Your 1 in 6 chance of rolling doubles now needs to be converted to probability. Probability speaks to the "chance" that something will occur. It is expressed as a "pure" number, and the range of probability is from 0 to 1. The probability 0 means that it cannot occur, and the probability 1 means that it will or must occur. Everything else falls in between. If you have a 1 in 6 chance of rolling doubles that translates into a 1 divided by 6 probability, or a 0.1666666... probability that a double will come up.
Why can't the probability of an event occurring be greater than 1?
If probability becomes greater than one, then the uncertainty coefficient fluctuates and results in a terminal case.
The answer would be 2/11 plus 1/2. Whatever that is.
If two fair dice are rolled what is the probability of rolling a double given that the sum was 11?
Let A = rolling a double Let B = sum is 11 P(A)=6/36=1/6 P(B)=2/36=1/18 since (5,6) and (6,5) produce a sum of 11. We want to find P(A/B)= P(A & B) / P(B) = 0 / P(B)=0 P(A & B) represent the event getting a double and the sum being 11.
A 50% to 50% chance (half to half).
Probability of heads in a coin toss?
1/2 Because there is one side with heads (1/) and the total sides (/2)
There are eight possible results when flipping three coins (eliminating the highly unlikely scenario of one or more coins landing on their edge): Dime - Heads / Nickel - Heads / Penny - Heads Dime - Heads / Nickel - Heads / Penny - Tails Dime - Heads / Nickel - Tails / Penny - Heads Dime - Heads / Nickel - Tails / Penny - Tails Dime - Tails / Nickel - Heads / Penny - Heads Dime - Tails / Nickel - Heads / Penny - Tails Dime - Tails / Nickel - Tails / Penny - Heads Dime - Tails / Nickel - Tails / Penny - Tails
What is the shape of this distribution?
A company is studying the number of daily debit card purchases. If there were 20 purchases and the probability of a debit card purchase is 0.5
What is the shape of this distribution
A three digit number can be chosen in 9*10*10 = 900ways
No. of ways to choose a three digit number without 7 is 8*9*9 = 648 ways
probability of picking a three digit number that includes atleast one digit that is 7 is 1- (648/900) = 252/900 = 7/25